# Seeing through Satellite Images at Street Views

### v1

@@ Abstract
This paper studies the task of SatStreet-view synthesis, which aims to render photorealistic street-view panorama images and videos given any satellite image and specified camera positions or trajectories. We formulate to learn neural radiance field from paired images captured from satellite and street viewpoints, which comes to be a challenging learning problem due to the sparse-view natural and the extremely-large viewpoint changes between satellite and street-view images. We tackle the challenges based on a task-specific observation that street-view specific elements, including the sky and illumination effects are only visible in street-view panoramas, and present a novel approach Sat2Density++ to accomplish the goal of photo-realistic street-view panoramas rendering by modeling these street-view specific in neural networks. In the experiments, our method is testified on both urban and suburban scene datasets, demonstrating that Sat2Density++ is capable of rendering photorealistic street-view panoramas that are consistent across multiple views and faithful to the satellite image. Project page is available at https://qianmingduowan.github.io/sat2density-pp/.
@@ 1 Introduction
Figure 1: Comparison of model designs and results between Sat2Density [1] proposed in our conference version and Sat2Density++ (Ours) proposed in this paper. (a) Overview illustration of model designs and the video results. At the bottom of Fig. 1(a), the colored trajectory of cameras are shown on the left satellite image and the corresponding street-view images generated by models are shown on the right. (b) A comparison of the generated geometry, where surfaces are extracted from the density field using Marching Cube. The satellite image is oriented with North at the top and West to the left. For the generated 360° panorama images, North aligns with the central axis, South spans the edge boundaries, while West and East occupy the left/right quarter sections respectively.
Please visit our project page for full video results.
@@
Reconstructing 3D scenes from multiple views of images is a core problem in the 3D vision and photogrammetry communities. This problem has been extensively studied in neural rendering [2, 3, 4], but usually suffers from a high demand of dense input views, often making it a resource-intensive practice, especially for outdoor scenes. In this paper, we focus on the 3D reconstruction of outdoor scenes in neural rendering, but aim to reduce the required number of input views to just two or even one using the wide-coverage characteristics of satellite images. We use a challenging two-view configuration that contains one satellite image capturing a large-scale scene from a distant location and one street view panorama capturing the same scene at street level, with the goal of feedforward learning faithful 3D scene geometry and rendering high-quality street-view videos at any queried bird’s-eye view (BEV) trajectory for the given satellite image.
@@
We call our two-view data as the SatStreet-view images, which pose an extremely challenging scenario for the establishment of visual correspondences and the neural rendering, but are viable by mutually and manually checking the context over the image pairs because both the satellite and street-view images are observing the same geographical contents.
Thus, we are curious about how we could infer the 3D scene geometry by curating the SatStreet-view image data. We take the novel-view synthesis as the main goal and refer to the task as the SatStreet-view synthesis.
@@
The SatStreet-view synthesis task aims to render photorealistic street-view panorama images and videos given any satellite image and specified camera positions or trajectories, as shown on Fig. 1. The primary challenge of this task stems from the significant viewpoint changes between satellite and street-level perspectives. Satellites capture scenes from a bird’s-eye view, mainly focusing on the tops of buildings, road surfaces, trees, and other geographical features. In contrast, the panoramas provide street-level views from within the scene, offering side perspectives of buildings, detailed road surfaces, the intricate features of tree trunks and branches, etc. Additionally, street-view panoramas include elements such as the sky and the view-specific illumination effects, which are beyond the limited scope of satellite imagery. To address these substantial viewpoint changes, we focus on the 3D scene geometry, which is likely invariant to observational perspectives.
@@
A robust 3D representation is essential for models to understand variations in camera viewpoints, thereby improving their ability to generate images more accurately. Shi et al. [5] proposed an end-to-end learning approach for height maps as an intermediate 3D geometric representation, demonstrating better results compared to image-to-image translation methods [6, 7] for single image generation. Other works [8, 9, 10] have circumvented the challenges of learning 3D scenes from SatStreet-view image data by employing off-the-shelf semantic and metric depth maps of satellite images as supervision signals, thereby obtaining 2.5D building voxels to serve as conditions for street-view images generation. All those studies highlighted the necessity of 3D shape representation for SatStreet-view synthesis, but they are limited by the accuracy and coherency of the used (or learned) 3D representations.
@@
Method Overview of Sat2Density++.  We improve the Sat2Density to Sat2Density++ in this paper. Compared to Sat2Density [1] that only learned the density field while ignoring color in the common neural field to reduce the complexity of the 3D representation, in our Sat2Density++, we demonstrate that both the scene geometry and the appearance modeling are playing the key roles to harness the complicated scenes, thus opt to learn an illumination-adaptive neural radiance field instead of the previously-used density field. This radiance field is capable of maintaining the characteristic hues of satellite imagery under neutral lighting conditions, or adjusting the generated color features when explicit illumination information is available. As a result, the density output of the illumination-adaptive neural radiance field is invariant across various lighting scenarios. Building atop of the joint modeling of scene geometry and the appearance, we simplify the generation of street-view panoramas in two parallel branches: one for the ground scene rendering from the learned neural radiance field and the other for the sky region generation from the illumination-sensitive 2D sky region generation network, and then compose the final street-view panoramas from the output of the two branches. By joint training of all the neural network components with carefully-designed learning objectives, we are able to learn the faithful 3D scene geometry and generate high-quality street-view panoramas that are consistent across multiple views and faithful to the satellite image, and thus can consecutively render street-view videos for any queried BEV trajectory. Fig. 1 illustrates the case for the urban scene modeling that are challenging the existing approaches (including our preliminary Sat2Density [1]) for Sat2Street-view synthesis, in which our proposed Sat2Density++ learns faithful 3D geometry and renders high fidelity street-view videos at any queried BEV trajectory.
@@ 2.1 Satellite-Ground Cross-view Perception
Both street-view and satellite images provide unique perspectives of the world, and their combination provides us with a more comprehensive way to understand and perceive the world from satellite-ground visual data. However, the drastic viewpoint changes between the satellite and ground images pose several challenges in geo-localization [11, 12, 13, 14, 15], SatStreet-view synthesis[5, 9, 8, 16, 10], cross-view synthesis [17, 7], overhead image segmentation with the assistance of ground-level images [18], geo-enabled depth estimation [19, 20], predicting ground-level scene layout from aerial imagery [21, 22].
To jointly extract more information from SatStreet-view images, many previous works have proposed various approaches to model and learn the drastic viewpoint changes, including the use of homography transforms [17], additional depth or semantic supervision [16, 8, 9], transformation matrices [11], and geospatial attention [18], among others. Despite effectiveness, most approaches mainly focus on modeling the feature correlations at the image level, while overlooking the intrinsic correspondence between the two views in 3D space. A seminal attempt by Shi et al. [5] utilizes height multi-plane images as the geometry representation on overcoming these challenges but fails to get good quality 3D representation from cross-view datasets. Our preliminary work [1] builds on this effort, revealing crucial insights that significant viewpoint variations and varying imaging conditions in cross-view images are the key factors hindering the learning of scene representation from cross-view images, and proposed approaches to aid in learning the scene representation.
@@ 2.2 Scene Generation from Conditions
In the urban context, innovations such as UrbanGIRAFFE [26] and InfiniCity [27] emphasize semantic voxel-based conditioning for scene creation. These 3D semantic voxels can comprehensively represent 3D semantics and occupancy information of the scene, reducing the learning burden on the model. However, annotating 3D semantic voxels requires extensive manual labor, which increases exponentially with the desire for more refined categories.
@@ 3 The Proposed Sat2Density++
Figure 2: Diagram of the proposed Sat2Density++ framework. The system begins with a satellite image input, it generates tri-plane features via the Tri-plane Net. Given specific camera poses, these features are then processed by an Illumination Adaptive Tri-plane Decoder within a Neural Renderer to render both satellite image and ground part of street view image. Additionally, a 2D sky generation module is responsible for creating the sky region in the street-view image. The final street-view images are obtained by first alpha-blending the ground and sky components, followed by super-resolution enhancement. The illumination input facilitates the rendering process by harmonizing both the Tri-plane Decoder and the 2D sky generation module. For clarity, we have omitted the steps involving the use of camera poses to generate image features from the radiance field and the super-resolution module, as well as the details of rendering from satellite viewpoints from the tri-plane.
This section introduces the details of our proposed Sat2Density++, as illustrated in Fig. 2.
@@ 3.1.1 Tri-plane Generation from Satellite Image
3.1.2 Street-View Illumination Modeling
As our goal is to learn how to synthesize street-view panoramas from satellite images, it is essential to model illumination conditions that are specific to street-level observations. Illumination is closely related to the sky region in street-view images $I_{\text{str}}$. Therefore, we adopt a non-learning approach to model the histogram-based illumination feature from each street-view image, which efficiently encodes color information while maintaining computational simplicity, interpretability, and does not rely on deep learning process.
@@ 3.1.2 Street-View Illumination Modeling
Remarks on the Inference Process. During inference, a street-view panorama does not exist prior to rendering, making it impossible to compute the illumination feature directly. However, we found in experiments that randomly sampling an illumination feature from the training set often produces plausible street-view panoramas. Furthermore, varying the sampled illumination feature leads to corresponding changes in the environmental lighting of the synthesized panorama. Therefore, in our default inference setting, we retain the use of precomputed color histograms from the training phase to guide the synthesis process.
@@ 3.1.3 Illumination-Adaptive Tri-Plane Decoder
Given the tri-plane feature $f_{\rm tri}$ for each sampled 3D location, we setup a illumination-adaptive tri-plane decoder for the computation of density value $\sigma\in\mathbb{R}$ and the appearance feature $\phi\in\mathbb{R}^{32}$ by
@@
where $\mathbf{w}_{\rm ill}\in\mathbb{R}^{512}$ is a style vector transformed from the illumination feature $f_{\rm ill}$ by an illumination mapping layer $E_{\rm ill}(\cdot)$ that is implemented by 8×8$\times$8 × fully connected layers, read as
@@
To input illumination feature, we inherit the MLP architecture used in the vanilla tri-plane decoder[28] but replace the single hidden layer with two parallel hidden layers. One hidden layer serves as the main rountine to produce the density value $\sigma$, and the other one concatenates the input of output feature of the main rountine with the illumination vector $\mathbf{w}_{\rm ill}$ as input to produce the apperance feature $\phi$. At inference time, $f_{\rm ill}$ can be set to any illumination input, facilitating illumination-controllable street-view video generation.
@@ 3.2 Street-View Image Generation
With the illumination-adaptive tri-plane representation, we focus on the generation of street-view panoramas by first leveraging volumetric rendering to yield the ground part and then generating sky images. An alpha blending of the rendered ground part and the generated sky part produces the final street-view panoramas.
@@ 3.2.1 Ground Part Generation
Cameras Definition. The panorama street-view camera utilizes a cylindrical projection model, positioning it horizontally with the camera’s compass heading aligned to true north. In the absence of height data for the street-view camera positions, we assume that all street images are captured at a uniform height above the ground.
@@
note that, the synthesis raw color $\hat{I}_{C}(r)$ is the first three layer of $\hat{I_{F}}(r)$, the other layers are used for super-resolution(SR) module.
Finally, given a street-view camera parameter, the model can generate ground part street-image feature $\hat{I}_{F\text{grd}}\in\mathbb{R}^{32\times H\times W}$, image color $\hat{I}_{C\text{grd}}\in\mathbb{R}^{3\times H\times W}$, opacity $\hat{O}_{\text{grd}}\in\mathbb{R}^{1\times H\times W}$, and depth $\hat{D}_{\text{grd}}\in\mathbb{R}^{1\times H\times W}$, the rendered size $(H,W)$ depends on the given camera intrinsic, in practice, the rendered size is $(64,256)$.
@@ 3.2.2 Sky Part Generation and Alpha Blending
Street-View Super Resolution. Similar to prior works [28, 32, 33, 34], as a final step, we apply a 2D image super-resolution (SR) module $U$ to reduce the computational cost for volume rendering when synthesizing street-view images. We get the high-resolution image $\hat{I}_{C\text{str}}^{+}$ by sending low-resolution $\hat{I}_{F\text{str}}$ to a lightweight SR module $U$:

### v2

@@ Abstract
This paper studies the task of SatStreet-view synthesis, which aims to render photorealistic street-view panorama images and videos given a satellite image and specified camera positions or trajectories. Our approach involves learning a satellite image conditioned neural radiance field from paired images captured from both satellite and street viewpoints, which comes to be a challenging learning problem due to the sparse-view nature and the extremely large viewpoint changes between satellite and street-view images. We tackle the challenges based on a task-specific observation that street-view specific elements, including the sky and illumination effects, are only visible in street-view panoramas, and present a novel approach, Sat2Density++, to accomplish the goal of photo-realistic street-view panorama rendering by modeling these street-view specific elements in neural networks. In the experiments, our method is evaluated on both urban and suburban scene datasets, demonstrating that Sat2Density++ is capable of rendering photorealistic street-view panoramas that are consistent across multiple views and faithful to the satellite image. Project page is available at https://qianmingduowan.github.io/sat2density-pp/.
@@ 1 Introduction
Figure 1: Comparison of model designs and results between Sat2Density [41] proposed in our conference version and Sat2Density++ (Ours) proposed in this paper. (a) Overview illustration of model designs and the video results. At the bottom of Fig. 1(a), the colored camera trajectory is shown on the left satellite image, and the corresponding street-view images generated by models are shown on the right. (b) A comparison of the generated geometry, where surfaces are extracted from the density field using Marching Cubes. Please visit our project page for full video results.
@@
Reconstructing 3D scenes from multiple views of images is a core problem in the 3D vision and photogrammetry communities. This problem has been extensively studied in neural rendering [37, 18, 44], but usually suffers from a high demand of dense input views, often making it a resource-intensive practice, especially for outdoor scenes. In this paper, we focus on the 3D reconstruction of outdoor scenes in neural rendering, but aim to reduce the required number of input views to just two or even one using the wide-coverage characteristics of satellite images. We use a challenging two-view configuration that contains one satellite image capturing a large-scale scene from a distant location and one street-view panorama capturing the same scene at street level, with the goal of learning a feedforward model that captures faithful 3D geometry and renders high-quality street-view videos at any queried bird’s-eye view (BEV) trajectory for the given satellite image.
@@
We refer to the two-view data as the SatStreet-view images, which pose an extremely challenging scenario for the establishment of visual correspondences and the neural rendering, but are viable by mutually and manually checking the context over the image pairs because both the satellite and street-view images are observing the same geographical contents. Thus, we are curious about how we could infer the 3D scene geometry by curating the SatStreet-view image data. We take the novel-view synthesis as the main goal and refer to the task as the SatStreet-view synthesis.
@@
The SatStreet-view synthesis task aims to render photorealistic street-view panorama images and videos given a satellite image and specified camera positions or trajectories, as shown in Fig. 1.  [^1] In this paper, all visualized satellite images are oriented with North at the top and West to the left. For the generated 360° panorama images, North aligns with the central axis, South spans the boundary edges, while West and East occupy the left and right quarter sections, respectively. The primary challenge of this task stems from the significant viewpoint changes between satellite and street-level perspectives. Satellites capture scenes from a bird’s-eye view, mainly focusing on the tops of buildings, road surfaces, trees, and other geographical features. In contrast, the panorama images provide street-level views from within the scene, offering side perspectives of buildings, detailed road surfaces, the intricate features of tree trunks and branches, etc. Additionally, street-view panorama images include elements such as the sky and the view-specific illumination effects, which are beyond the limited scope of satellite imagery. To address these substantial viewpoint changes, we focus on the 3D scene geometry, which is likely invariant to observational perspectives.
@@
A robust 3D representation is essential for models to understand variations in camera viewpoints, thereby improving their ability to generate images more accurately. Shi et al. [46] proposed an end-to-end learning approach for height maps as an intermediate 3D geometric representation, demonstrating better results compared to image-to-image translation methods [15, 42] for single image generation. Other works [34, 30, 29] have circumvented the challenges of learning 3D scenes from SatStreet-view image data by employing off-the-shelf semantic and metric depth maps of satellite images as supervision signals, thereby obtaining 2.5D building voxels to serve as conditions for street-view images generation. All those studies highlighted the necessity of 3D shape representation for SatStreet-view synthesis, but they are limited by the accuracy and coherence of the used (or learned) 3D representations.
@@
Method Overview of Sat2Density++.  We extend the Sat2Density to Sat2Density++ in this paper. Compared to Sat2Density [41] that only learned the density field while ignoring color in the common neural field to reduce the complexity of the 3D representation, in our Sat2Density++, we demonstrate that both the scene geometry and the appearance modeling play key roles in modeling complex scenes, thus opt to learn an illumination-adaptive neural radiance field instead of the previously-used density field. This radiance field is capable of maintaining the characteristic hues of satellite imagery under neutral lighting conditions or adjusting the generated color features when explicit illumination information is available. As a result, the density output of the illumination-adaptive neural radiance field is invariant across various lighting scenarios. Building on the joint modeling of scene geometry and the appearance, we simplify the generation of street-view panoramas in two parallel branches: one for the ground scene rendering from the learned neural radiance field and the other for the sky region generation from the illumination-sensitive 2D sky region generation network, and then compose the final street-view panoramas from the output of the two branches. By joint training of all the neural network components with carefully-designed learning objectives, we are able to learn the faithful 3D scene geometry and generate high-quality street-view panoramas that are consistent across multiple views and faithful to the satellite image, and thus can consecutively render street-view videos for any queried BEV trajectory. Fig. 1 illustrates the case for the urban scene modeling that is challenging the existing approaches (including our preliminary Sat2Density [41]) for SatStreet-view synthesis, in which our proposed Sat2Density++ learns faithful 3D geometry and renders high-fidelity street-view videos at any queried BEV trajectory.
@@ 2.1 Satellite-Ground Cross-view Perception
Both street-view and satellite images provide unique perspectives of the world, and their combination provides us with a more comprehensive way to understand and perceive the world from satellite-ground visual data. However, the drastic viewpoint changes between the satellite and ground images pose several challenges in geo-localization [67, 48, 49, 51, 25, 16, 53, 60, 47, 7, 45, 50, 61, 52], SatStreet-view synthesis[46, 30, 34, 54, 29, 27, 66], cross-view synthesis [43, 42, 32, 65], street-view image synthesis from multiple satellite images [63, 8], overhead image segmentation with the assistance of ground-level images [59], geo-enabled depth estimation [58, 13], and predicting ground-level scene layout from aerial imagery [68, 26].
To jointly extract more information from SatStreet-view images, many previous works have proposed various approaches to model and learn the drastic viewpoint changes, including the use of homography transforms [43], additional depth or semantic supervision [54, 34, 30], transformation matrices [67], and geospatial attention [59], among others. Despite effectiveness, most approaches mainly focus on modeling the feature correlations at the image level, while overlooking the intrinsic correspondence between the two views in 3D space. A seminal attempt by Shi et al. [46] utilizes height multi-plane images as the geometry representation to overcome these challenges, but fails to get a good quality 3D representation from cross-view datasets. Our preliminary work [41] builds on this effort, revealing crucial insights that significant viewpoint variations and varying imaging conditions in cross-view images are the key factors hindering the learning of scene representation from cross-view images, and proposes approaches to aid in learning the scene representation.
@@ 2.2 Scene Generation from Conditions
In the urban context, innovations such as UrbanGIRAFFE [64], Streetscapes [5], and InfiniCity [31] emphasize semantic voxel-based conditioning for scene creation. These 3D semantic voxels can comprehensively represent 3D semantics and occupancy information of the scene, reducing the learning burden on the model. However, annotating 3D semantic voxels requires extensive manual labor, which increases exponentially with the desire for more refined categories.
@@ 3 The Proposed Sat2Density++
Figure 2: Diagram of the proposed Sat2Density++ framework. The system begins with a satellite image input, and it generates tri-plane features via the Tri-plane Net. Given specific camera poses, these features are then processed by an Illumination Adaptive Tri-plane Decoder within a Neural Renderer to render both the satellite image and the ground part of the street-view image. Additionally, a 2D sky generation module is responsible for creating the sky region in the street-view image. The final street-view images are obtained by first alpha-blending the ground and sky components, followed by super-resolution enhancement. The illumination input facilitates the rendering process by harmonizing both the Tri-plane Decoder and the 2D sky generation module. For clarity, we have omitted the steps involving the use of camera poses to generate image features from the radiance field and the super-resolution module, as well as the details of rendering from satellite viewpoints from the tri-plane.
This section introduces the details of our proposed Sat2Density++, as illustrated in Fig. 2. In our model, the roles of the satellite image and the sky illumination input are completely decoupled. The satellite image is responsible for generating a learnable implicit 3D representation of outdoor ground scenes. Using the given camera parameters for the satellite or panoramic street view, it can accurately produce the corresponding ground content. In contrast, the infinite sky region is treated as a 2D background panoramic image, which is generated using the sky illumination input. The sky illumination input simulates the effects of varying daylight conditions on photographs in outdoor environments, with different daylight scenarios altering both the sky appearance and impacting the imaging of ground regions. Consequently, in our model, the sky illumination input also influences the appearance of images when rendering the street region from the implicit 3D representation, allowing for dynamic adaptation to different lighting conditions.
@@ 3.1.1 Tri-plane Generation from Satellite Image
3.1.2 Sky Illumination Modeling from Street-View Image
As our goal is to learn how to synthesize street-view panoramas from satellite images, it is essential to model illumination conditions that are specific to street-level observations. Illumination is closely related to the sky region in street-view images $I_{\text{str}}$. Therefore, we adopt a statistical approach to model the histogram-based illumination feature from each street-view image, which efficiently encodes color information while maintaining computational simplicity, interpretability, and does not rely on a deep learning process. Lalonde et al. [21] also utilized sky histograms to match global lighting conditions.
@@ 3.1.2 Sky Illumination Modeling from Street-View Image
Remarks on the Inference Process. Leveraging our approach, which decouples the sky illumination module from 3D implicit scene generation, we are able to produce street-view images or videos under various daylight conditions from a single satellite image by applying different sky illumination inputs. In our quantitative tests, we adopted multiple strategies. For instance, results marked with ‘*’ in Tab. I illustrate that each satellite image was paired with a randomly selected sky illumination from the training set for rendering street views. Conversely, results marked with ‘†’ in Tab. I utilized sky illumination inputs derived from the corresponding ground truth street-view images. In our qualitative results, we further showcased the capability to generate street views under different lighting conditions while maintaining a constant satellite image, as depicted in Fig. 9 and Fig. 12.
@@ 3.1.3 Illumination-Adaptive Tri-Plane Decoder
Given the tri-plane feature $f_{\rm tri}$ for each sampled 3D location, we set up an illumination-adaptive tri-plane decoder for the computation of density value $\sigma\in\mathbb{R}$ and the appearance feature $\phi\in\mathbb{R}^{32}$ by
@@
where $\mathbf{w}_{\rm ill}\in\mathbb{R}^{512}$ is a style vector transformed from the illumination feature $f_{\rm ill}$ by an illumination mapping layer $E_{\rm ill}(\cdot)$ that is implemented by $8\times$ fully connected layers, read as
@@
The vanilla tri-plane decoder consists of a two-layer structure formed by multilayer perceptrons (MLPs). In this model, $f_{\rm tri}$ are processed through these MLPs, producing $\phi$ and $\sigma$ outputs simultaneously in the final layer. This approach assumes consistent lighting conditions across all viewpoints, providing a straightforward output based solely on triplane features [2].
In contrast, the illumination adaptive tri-plane decoder modifies this structure to address varying lighting conditions more effectively. Instead of using two stacked MLP layers as in the vanilla version, the illumination adaptive version replaces the final layer with two distinct branches for generating density and color, respectively. The density branch utilizes an MLP that takes the output from the first MLP layer as input to produce $\sigma$. The color branch also uses an MLP, but it concatenates the output from the first MLP layer with the illumination vector $\mathbf{w}_{\rm ill}$ before processing to generate $\phi$. This design allows the color features to be adaptively adjusted according to lighting conditions, offering more nuanced control and improving the accuracy and realism of visual representation under different illumination contexts.
@@ 3.2 Street-View Image Generation
With the illumination-adaptive tri-plane representation, we focus on the generation of street-view panoramas by first leveraging volumetric rendering to yield the ground part and then generating a sky image. An alpha blending of the rendered ground part and the generated sky part produces the final street-view panoramas.
@@ 3.2.1 Ground Part Generation
Camera Definition. The panorama street-view camera utilizes a cylindrical projection model, positioning it horizontally with the camera’s compass heading aligned to true north. In the absence of height data for the street-view camera positions, we assume that all street images are captured at a uniform height above the ground.
@@
note that, the synthesized raw color $\hat{I}_{C}(r)$ is taken from the first three channels of $\hat{I_{F}}(r)$, the other channels are used for the super-resolution (SR) module.
Finally, given a street-view camera parameter, the model can generate ground part street-image feature $\hat{I}_{F\text{grd}}\in\mathbb{R}^{32\times H\times W}$, image color $\hat{I}_{C\text{grd}}\in\mathbb{R}^{3\times H\times W}$, opacity $\hat{O}_{\text{grd}}\in\mathbb{R}^{1\times H\times W}$, and depth $\hat{D}_{\text{grd}}\in\mathbb{R}^{1\times H\times W}$, the rendered size $(H,W)$ depends on the given camera intrinsics, in practice, the rendered size is $(64,256)$.
@@ 3.2.2 Sky Part Generation and Alpha Blending
Street-View Super Resolution. Similar to prior works [2, 6, 10, 39], as a final step, we apply a 2D image super-resolution (SR) module $U$ to reduce the computational cost for volume rendering when synthesizing street-view images. We obtain the high-resolution image $\hat{I}_{C\text{str}}^{+}$ by feeding low-resolution $\hat{I}_{F\text{str}}$ into a lightweight SR module $U$: